Solve Problems Using Linear Systems
6. Consider the linear systems
3x-y=8
4x-y=-15
a) Which method would you choose to solve the linear systems and why? Solve using the method you chose.
b) Now solve using one of the other methods available to you.
(The answer is: a)(-23, -77) and b) (-23,-77), but i dont know how to solve it?) Answer by checkley77(12844) 
 3x-y=8 multiply this equation by -1 & add them.
4x-y=-15
-3x+y=-8
------------------------
x=-23 ans.
3*-23-y=8
-69-y=8
-y=8+69
-y=77
y=-77 ans.
Proof:
4*-23-(-77)=-15
-92+77=-15
-15=-15